15. Calculate the total number of bits transferred if 200 pages of ASCIIl data are sent using asynchronous serial data transfer. Assume a data size of 8 bits, I stop bit, and no parity. Assume each page has 80x25 of text characters. 16. In Problem 15, how long will the data transfer take if the baud rate is 9.600?

Please explain!
Answer #1

Answer:

For one character 8 bits are necessary and 1 stub bit

For one page 80X25 characters are present

Means for one page 80*25 = 2000 characters

For one page

Total no of bits = 2000*8 +2000

=18000 bits

For 200 pages

Total no of bits = 18000*200

=3600000 bits are transferred